# Understanding the temperature response of biological systems: Part I - Phenomenological descriptions and microscopic models

### v2

@@ Abstract
Virtually every biological rate depends on temperature, yet the resulting rate–temperature relationships often deviate strongly from simple Arrhenius behavior. In this first part of a two-part review, we survey empirical and phenomenological models used to describe biological temperature responses across scales, from enzymatic reactions to organismal performance. We discuss common functional forms, including symmetric and asymmetric thermal performance curves and extensions of the Arrhenius law, and we highlight how these models define operational quantities such as optimal temperatures, thermal breadths, and thermal limits. In Part II of this review, we will discuss how system-level temperature response curves emerge from the interaction of many underlying reactions.
@@ Introduction
Temperature sets the pace of life by modulating molecular stability, diffusion constants, reaction rates, and material properties. From enzymatic catalysis to developmental timing, virtually every biological process depends on temperature [1, 2, 3, 4, 5]. However, our understanding of the relevant mechanisms is not complete, which makes it difficult to understand how temperature shapes biological function across scales, ranging from the motion of individual molecules to the physiology of organisms and the behavior of ecosystems (Fig. 1A) [6, 7, 8, 9, 10].
@@
Because virtually every biologically relevant process depends on temperature, even modest temperature changes can have strong physiological effects. Many biological systems have therefore evolved mechanisms of temperature compensation, such as circadian clocks [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and bacterial chemotaxis [23, 24], that maintain function across environmental fluctuations. Other systems exploit temperature variations to effect a desired functional response, as in fever-mediated immune responses [25] or in temperature-dependent sex determination [26, 27, 28, 29, 30, 31, 32]. Many species employing the latter strategy, such as many reptiles and some fish, are increasingly threatened by climate change [33, 34, 35, 36]. These evolutionary adaptations prove the importance of temperature effects in biological systems, but in order to understand how they work, one needs to develop a mechanistic understanding of the effect of temperature on biological systems.
@@
The most fundamental effect of temperature on biological systems is that it determines the level of kinetic energy of the biomolecules. This is captured by the Stokes–Einstein relation $D=k_{B}T/6\pi\eta R$, which expresses the translational diffusion constant $D$ as a function of temperature $T$, viscosity $\eta$ and molecular radius $R$. The Boltzmann constant is denoted as $k_{B}$. For biomolecules with nanometer-scale dimensions in aqueous solution at body temperature, $D$ is typically on the order of $\mu m^{2}/s$. Experimental studies in E. coli have confirmed the fundamental validity of the Stokes–Einstein relation for biological systems [37]. Because biological cells work at relatively high temperatures, their diffusion constant are large and lead to frequent encounters between the biomolecules, which is essential for biochemical function and ensure that biological systems can quickly respond to environmental changes [38, 39].
@@
Figure 1: Temperature influences biological systems across scales and can be described using different classes of models. (A) Temperature acts from the level of particles and biochemical reactions to organisms and ecosystems. (B) Examples of empirical rate–temperature relationships: growth rates of a M. aeruginosa cyanobacteria colony [52] and cleavage rates during early development of D. rerio zebrafish embryos [53]. Both exhibit strong, nonlinear temperature dependence. (C) Conceptual illustration of Arrhenius and non-Arrhenius behavior. In an Arrhenius plot (log rate versus $1/T$), simple reactions follow a straight line, whereas biological processes typically show curvature, an optimum temperature $T_{\mathrm{o}}$, a maximal rate $r_{\mathrm{o}}$, and thermal limits ($T_{\min},T_{\max}$). (D) Overview of modeling frameworks used to describe temperature responses, organized by level of description. Phenomenological models provide empirical fits to observed rate–temperature curves, while microscopic models derive rate–temperature relationships from reaction-level kinetics. At a higher level, network-level models—either deterministic or stochastic—capture how temperature affects coupled biochemical or regulatory systems. The phenomenological and microscopic approaches form the focus of Part I [54], whereas deterministic and stochastic network models are the focus of Part II [55]. All four approaches are illustrated by fitting the same zebrafish cleavage-timing dataset [53], demonstrating how distinct model classes can reproduce the characteristic non-Arrhenius shape of biological temperature–response curves.
@@
Several reviews have compared empirical fitting functions and discussed their statistical performance and biological interpretation [59, 60, 46, 45]. Here, we take a different approach by organizing models according to their level of description, as summarized in Fig. 1D. In this first part, we focus on phenomenological models, which provide compact empirical descriptions of rate–temperature curves and thermal performance curves (TPCs) and define operational quantities such as $T_{\mathrm{o}}$, $W$, and thermal limits. We then review microscopic single-reaction level theories—including Arrhenius, Eyring, Kramers, and enzyme-stability models—which derive temperature dependence from physical and chemical principles but treat reactions in isolation. Together, these approaches describe how temperature affects individual processes and observed rate curves, while remaining agnostic about how system-level temperature responses emerge from interacting pathways. Network-level deterministic and stochastic models that address this emergence are discussed in Part II [55].
@@
[…]
Figure 2: Phenomenological and microscopic models of temperature responses. (Top) Common phenomenological approaches for fitting rate–temperature curves, including symmetric models, asymmetric models with distinct cold and warm widths, and extensions of the Arrhenius law that incorporate optimal temperatures and upper/lower limits. (Bottom) Schematic overview of key microscopic reaction–level theories: transition-state formulations such as Arrhenius and Eyring equations, Kramers’ barrier-crossing dynamics, and enzyme-catalyzed reactions with temperature-dependent active fractions.
@@
Mitchell-Angilletta model. Lastly, [46] found that a symmetric cosinuisodal symmetric temperature response curve, originally proposed by Mitchell and Angilletta as a simple mathematical model for the thermal adaptation of lizards [67], scores high on the AIC for experimental data across different traits and species, meaning that it combines a good fit with a small number of free parameters and mathematical simplicity. This curve has symmetric thermal limits, but in contrast to the quadratic polynomial, it approaches the extremal temperatures smoothly.
@@
Linearly varying $Q_{10}$. In the case of ideal Arrhenius scaling, the $Q_{10}$ factor of a process, which gives the increase in rate over a range of $10$ degrees, is constant. Atkin [76] therefore proposed to absorb deviations into a linearly temperature-dependent $Q_{10}$. This results in an asymmetric curve with interpretable limits that is easy to fit to experimental datasets [46].
@@
Microscopic models aim to derive rate–temperature relationships directly from the physics and chemistry of elementary reactions. They treat biological processes as thermally activated transitions, whether barrier crossing, formation of a transition state, or enzyme-mediated catalysis. In doing so they provide a mechanistic foundation for the exponential rise of rates with temperature and the deviations from Arrhenius scaling observed near thermal limits. In contrast to phenomenological fits, which summarize the shape of empirical curves, microscopic models address the origin of temperature dependence at the level of reaction steps. Below, we outline the central theoretical frameworks developed for this purpose.
@@ Acknowledgements
The work is supported by grants from Internal funds KU Leuven (C14/23/130, LG). JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support.

### v3

@@ Abstract
Virtually every biological rate depends on temperature, yet the resulting rate–temperature relationships often deviate strongly from simple Arrhenius behavior. In this first part of a two-part review, we survey phenomenological models used to describe biological temperature responses across scales, from enzymatic reactions to organismal performance. We discuss common functional forms, including symmetric and asymmetric thermal performance curves and extensions of the Arrhenius law, and we highlight how these models define operational quantities such as optimal temperatures, thermal breadths, and thermal limits. We also discuss microscopic models for the effect of temperature, which however do not capture cooperative effects. In Part II of this review, we will discuss how system-level temperature response curves emerge from the interaction of many underlying reactions.
@@ Introduction
Temperature sets the pace of life by modulating molecular stability, diffusion constants, reaction rates, and material properties. From enzymatic catalysis to developmental timing, virtually every biological process depends on temperature [1, 2, 3, 4, 5]. However, our understanding of the relevant mechanisms is not complete, which makes it difficult to understand how temperature shapes biological function across scales, ranging from the motion of individual molecules and the performance of biochemical reactions and their networks to the physiology of organisms and the behavior of ecosystems (Fig. 1A) [6, 7, 8, 9, 10].
@@
Because virtually every biologically relevant process depends on temperature, even modest temperature changes can have strong physiological effects. Many biological systems, guided by the biochemical adaptation strategies [11], have therefore evolved mechanisms of temperature compensation, such as circadian clocks [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and bacterial chemotaxis [25, 26], that maintain function across environmental fluctuations. Other systems exploit temperature variations to effect a desired functional response, as in fever-mediated immune responses [27] or in temperature-dependent sex determination [28, 29, 30, 31, 32, 33, 34]. Many species employing the latter strategy, such as many reptiles and some fish, are increasingly threatened by climate change [35, 36, 37, 38]. These evolutionary adaptations prove the importance of temperature effects in biological systems, but in order to understand how they work, one needs to develop a mechanistic understanding of the effect of temperature on biological systems.
From the materials point of view, biological systems are complex fluids with an aqueous solvent and thus their temperature $T$ is defined like for any thermodynamic system, namely as conjugate quantity to entropy $S$. Therefore energy in the form of heat is written as $dE=TdS$ and temperature is defined by $1/T=dS/dE$, with other state variables (such as particle number $N$ and volume $V$) held constant [39]. This definition makes sure that two systems at equilibrium will attain the same temperature $T$, and that out of equilibrium heat will flow from the warm to the cold system. For an ideal gas, for which each of the $N$ particles carries only kinetic energy $E=mv^{2}/2$, one can calculate the energy-dependent part of the entropy to be $S=Nk_{B}\ln E^{3/2}$, where $k_{B}$ is the Boltzmann constant. The factor $3$ arises because we live in a three-dimensional world, while the factor $2$ arises from the quadratic energy relation. Calculating the derivative $dS/dE$, setting it equal to $1/T$ and solving for $T$, one gets $E=3Nk_{B}T/2$. This result is an example of the equipartition theorem from statistical physics, which states that each degree of freedom ($3N$ for the ideal gas, where each particle can move in three directions) carries half the thermal energy $k_{B}T$. For biological systems at room or body temperature, thermal energy is $k_{B}T=4.1\ 10^{-21}$ J. In computer simulations of biological systems, a certain value of temperature $T$ usually is fixed by assuming that it is connected to a heat reservoir and by adjusting kinetic energies such that velocity $v=\sqrt{k_{B}T/m}$ for each direction of space [40]. In this review, we assume that temperature has exactly this thermodynamic role of fixing the energy level in the system.
Figure 1: Temperature influences biological systems across scales and can be described using different classes of models. (A) Temperature acts from the level of biomolecules and biochemical reactions and their networks to organisms and ecosystems. (B) Examples of empirical rate–temperature relationships: growth rates of a M. aeruginosa cyanobacteria colony [41] and cleavage rates during early development of D. rerio zebrafish embryos [42]. Both exhibit strong, nonlinear temperature dependence. (C) Conceptual illustration of Arrhenius and non-Arrhenius behavior. In an Arrhenius plot (log rate versus $1/T$), simple reactions follow a straight line, whereas biological processes typically show curvature, an optimum temperature $T_{\mathrm{o}}$, a maximal rate $r_{\mathrm{o}}$, and thermal limits ($T_{\min},T_{\max}$). (D) (Left) Overview of modeling frameworks used to describe temperature responses, organized by level of description. Phenomenological models provide empirical fits to observed rate–temperature curves, while microscopic models derive rate–temperature relationships from reaction-level kinetics. At a higher level, network-level models—either deterministic or stochastic—capture how temperature affects coupled biochemical or regulatory systems. The phenomenological and microscopic approaches form the focus of Part I [43], whereas deterministic and stochastic network models are the focus of Part II [44]. (Right) All four approaches are illustrated by fitting the same zebrafish cleavage-timing dataset [42], demonstrating how distinct model classes can reproduce the characteristic non-Arrhenius shape of biological temperature–response curves.
@@
Because temperature determines the level of kinetic energy of the biomolecules, it also determines their diffusion constant, which arises from the frequent collisions with solvent molecules and other biomolecules. This is captured by the Einstein relation $D=\mu k_{B}T$, where $\mu$ is the mobility and usually taken to be $\mu=1/6\pi\eta R$ as derived from the Stokes equation, with $\eta$ viscosity and $R$ the hydrodynamic radius. For biomolecules with nanometer-scale dimensions in aqueous solution at body temperature, $D$ is typically on the order of $\mu m^{2}/s$. Experimental studies in E. coli have confirmed the fundamental validity of the Stokes–Einstein relation for biological systems [45]. Because biological cells work at relatively high temperatures, their diffusion constants are large and lead to frequent encounters between the biomolecules, which is essential for biochemical function and ensures that biological systems can quickly respond to environmental changes [46, 47]. If non-equilibrium processes increase the energy level in the system, the Einstein relation can also be used to define an effective temperature $T_{eff}=D/\mu k_{B}$ . Such effects have been demonstrated for single molecules [48], the cytoskeleton [49] and the plasma membrane [50], but here we only discuss cases in which the equilibrium definition of temperature applies.
@@
Several reviews have compared empirical fitting functions and discussed their statistical performance and biological interpretation [66, 67, 57, 56]. Here, we take a different approach by organizing models according to their level of description, as summarized in Fig. 1D. In this first part, we focus on phenomenological models, which provide compact empirical descriptions of rate–temperature curves and thermal performance curves and define operational quantities such as $T_{\mathrm{o}}$, $W$, and thermal limits. We then review microscopic single-reaction level theories—including Arrhenius, Eyring, Kramers, and enzyme-stability models—which derive temperature dependence from physical and chemical principles but treat reactions in isolation. Together, these approaches describe how temperature affects individual processes and observed rate curves, while remaining agnostic about how system-level temperature responses emerge from interacting pathways. Network-level deterministic and stochastic models that address this emergence are discussed in Part II [44].
@@
Because many functional forms have been proposed, model choice requires balancing flexibility, interpretability, and parsimony. Information–theoretic analyses such as the Akaike information criterium (AIC) [72, 69, 57] show that simple three–parameter models often perform as well as or better than more complicated alternatives, and that no single functional form is universally optimal. Below, we organize commonly used three- or four-parameter phenomenological frameworks into symmetric models, asymmetric models, and Arrhenius–based extensions, highlighting those that perform well across diverse thermal datasets. In Supplemental Table I and II we have collected their explicit functional forms, which we re-parametrized in terms of $r_{\text{o}}$, $T_{\text{o}}$, $W$, $T_{\text{min}}$ and $T_{\text{max}}$ where possible. That such a comparison is possible, highlights the universality across phenomenological models.
Figure 2: Phenomenological and microscopic models of temperature responses. (Top) Common phenomenological approaches for fitting rate–temperature curves, including symmetric models, asymmetric models with distinct cold and warm widths, and extensions of the Arrhenius law that incorporate optimal temperatures and upper/lower limits. (Bottom) Schematic overview of key microscopic reaction–level theories: transition-state formulations such as Arrhenius and Eyring equations (with parameters coded by respective colors), Kramers’ barrier-crossing dynamics, and enzyme-catalyzed reactions with temperature-dependent active fractions. Color-coding of phenomenological and microscopic models refers to Figure 1D.
@@
Mitchell-Angilletta model. Lastly, [57] found that a symmetric cosinusodal symmetric temperature response curve, originally proposed by Mitchell and Angilletta as a simple mathematical model for the thermal adaptation of lizards [74], scores high on the AIC for experimental data across different traits and species, meaning that it combines a good fit with a small number of free parameters and mathematical simplicity. This curve has symmetric thermal limits, but in contrast to the quadratic polynomial, it approaches the extremal temperatures smoothly.
@@
Varying $Q_{10}$. In the case of ideal Arrhenius scaling, the $Q_{10}$ factor of a process, which gives the increase in rate over a range of $10$ degrees, is constant. While typical values for chemical reactions are $Q_{10}=2-3$, a value greater than unity cannot account for the decline in performance at the hot side of a biological TPC [13, 83]. Therefore various authors have suggested to fit TPCs with a $Q_{1}0$ value that declines with temperature. [84, 85]. A particularly simple proposal is that of Atkin [86] who came up with a linearly temperature-dependent $Q_{10}$. This results in an asymmetric curve with interpretable limits that is easy to fit to experimental datasets [57].
@@
Microscopic models aim to derive rate–temperature relationships directly from the physics and chemistry of elementary reactions. They treat biological processes as thermally activated transitions, whether barrier crossing, formation of a transition state, or enzyme-mediated catalysis. In doing so they provide a mechanistic foundation for the exponential rise of rates with temperature and the deviations from Arrhenius scaling observed near thermal limits. In contrast to phenomenological fits, which summarize the shape of empirical curves, microscopic models address the origin of temperature dependence at the level of reaction steps. Below, we outline the central theoretical frameworks developed for this purpose. For a more comprehensive overview of the fundamental principles of chemical and enzyme kinetics, we refer the readers to Chapters I and II of a seminal textbook by Cornish-Bowden [89].
@@ Conclusion
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used ChatGPT in order to get feedback on grammar and phrasing. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Acknowledgements
The work is supported by grants from Internal funds KU Leuven (C14/23/130, LG). JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) and USS the Max Planck School Matter to Life, which is funded by the Dieter Schwarz Foundation and the Max Planck Society.
Table of acronyms and technical terms
Term Meaning
AIC Akaike information criterium
EAAR enzyme-assited Arrhenius
TPC thermal performance curve
TST transition-state theory
Table 1: Alphabetical table of acronyms and technical terms used in the main text and their meaning.